# A Probabilistic Parking Process and Labeled IDLA

### v1

@@ 5.1. Consequences of Negative Correlation
$\begin{split}\operatorname{Pr}\left[N_{\alpha}\geq(1+\delta)\sum\limits_{i=1}^{n}\Pr\left[X_{\alpha}^{i}=1\right]\right]&\leq\left(\frac{e^{\delta}}{(1+\delta)^{1+\delta}}\right)^{\sum\limits_{i=1}^{n}\Pr\left[X_{\alpha}^{i}=1\right]}\\ \end{split}$
@@
$\begin{split}\operatorname{Pr}\left[N_{\alpha}\leq(1-\delta)\sum\limits_{i=1}^{n}\Pr\left[X_{\alpha}^{i}=1\right]\right]&\leq\left(\frac{e^{\delta}}{(1-\delta)^{1-\delta}}\right)^{\sum\limits_{i=1}^{n}\Pr\left[X_{\alpha}^{i}=1\right]}\,.\\ \end{split}$
@@ 6. Future work and Open Questions
  • Statistics of the Outcome. In general, it would be interesting to better understand statistics of the outcome of the probabilistic parking process for general $\alpha\in[n]^{n}$. For example, we know that, if we remove the assumption that $\alpha\in\mathrm{PF}_{n}(\mathrm{id})$, at the end of the probabilistic parking process the parking lot could possibly have “holes,” or spots where no car has parked. Suppose that we have run the probabilistic parking protocol with preference list $\alpha\in[n]^{n}$, and define $B_{\alpha}$ to be a tuple of length $n$ where $B_{\alpha}(i)=1$ if spot $i$ is filled at the end of the parking process and 00 otherwise. What is the probability that $B_{\alpha}=b$ for some $b\in\{0,1\}^{n}$? This line of research is closely related to the work of Varin, and thus the techniques in [43] might be useful. However, because we also insist on distinguishing between the cars throughout the process, one could study the probability that a specific car, say car $i$, parks in spot $j$ for $i,j\in[n]$, with or without the assumption that all of the cars are able to successfully park. If all the cars successfully park, we could ask what is the probability that a specific permutation $\pi\in S_{n}$ is the outcome of the parking process started with a preference list $\alpha\in[n]^{n}$.
@@
  • Other Probabilistic Parking Protocols. As mentioned in the introduction, parking has been considered on random objects such as random trees [23] and frozen Erdős Renyi graphs [11], as far as we are aware, the present paper and the paper of Durmić et al. [18], the paper of Tian and Trevino [42], and the paper of Varin [43] are the only works which consider randomizing the parking protocol itself. As discussed, the model introduced in this paper has deep connections to the IDLA particle model, and thus it seems interesting to derive a new probabilistic parking protocol which relates to (or perhaps is independent of) other more classical particle models or stochastic processes.
@@ Appendix A Supplementary Results
$\hskip 36.135pt+n\binom{2n-1}{n-2}$
[…]
$\hskip 36.135pt+n\binom{2n-1}{n-2}$
$=\frac{n^{3}(n-1)}{(n+1)(n+2)}\binom{2n-1}{n-1}.\qed$
@@
Since the elements in $A_{n}(j)$ are uniquely determined by their first $n-1$ coordinates, and since any prefix belonging to both $B_{n}(j)$ and $B_{n}(j-t)$ for $12\leq t\leq j-$ 1 also belongs to $B_{n}(j-1)$, we conclude
@@
The case of $g_{0}=g_{i}=0$ is easily shown to be satisfied. For $11\leq s\leq i-$ 1, use that
@@
(A.4) $\frac{\left(1+\left(\frac{q}{p}\right)^{s+1}\right)}{\left(1-\left(\frac{q}{p}\right)^{s+1}\right)}=\frac{p\cdot p^{s}+q\cdot q^{s}}{p\cdot p^{s}-q\cdot q^{s}}\ \ \mbox{ and }\ \ \frac{\left(1+\left(\frac{q}{p}\right)^{s-1}\right)}{\left(1-\left(\frac{q}{p}\right)^{s-1}\right)}=\frac{q\cdot p^{s}+p\cdot q^{s}}{q\cdot p^{s}-p\cdot q^{s}}\,.$
Next we evaluate the last case in (A.2) using (A.1). By (A.3), $p\frac{w_{s+1}}{w}+q\frac{w_{s-1}}{w}=1$ and so our only concern with respect to each term on the right hand side of (A.1), is the $-\frac{s\left(1+\left(\frac{q}{p}\right)^{s\pm 1}\right)}{(p-q)\left(1-\left(\frac{q}{p}\right)^{s\pm 1}\right)}$ contribution. Thus, using (A.3) and (A.4), we have that
@@ Acknowledgments
The authors gratefully acknowledge that this collaboration began at the 2024 Graduate Research Workshop in Combinatorics, which is supported in part by NSF Grant DMS – 1953445. The authors would like to sincerely thank Steve Butler for initial contributions and providing a key insight in Proposition 2.2. The authors also thank Alice Contat and Ivailo Hartarsky for fruitful discussion, and for introducing us to similar models in the literature. Finally, we would like to thank Persi Diaconis for taking the time to speak with us about this project, for helpful suggestions, and for offering his unique perspective on the model. Parts of this project were also presented at the Oberwolfach Mini-Workshop: Mixing Times in the Kardar-Parisi-Zhang Universality Class where useful discoveries were made, and helpful discussions were had.

### v2

@@ 5.1. Consequences of Negative Correlation
[…]
@@
[…]
@@ 6. Future work and Open Questions
  • Statistics of the Outcome. In general, it would be interesting to better understand statistics of the outcome of the probabilistic parking process for general $\alpha\in[n]^{n}$. For example, we know that, if we remove the assumption that $\alpha\in\mathrm{PF}_{n}(\mathrm{id})$, at the end of the probabilistic parking process the parking lot could possibly have “holes,” or spots where no car has parked. Suppose that we have run the probabilistic parking protocol with preference list $\alpha\in[n]^{n}$, and define $B^{\alpha}$ to be a tuple of length $n$ where $B^{\alpha}_{i}=1$ if spot $i$ is filled at the end of the parking process and 0 otherwise. What is the probability that $B^{\alpha}=\beta$ for some $\beta\in\{0,1\}^{n}$? This line of research is closely related to the work of Varin, and thus the techniques in [VarinGolf] might be useful. However, because we also insist on distinguishing between the cars throughout the process, one could study the probability that a specific car, say car $i$, parks in spot $j$ for $i,j\in[n]$, with or without the assumption that all of the cars are able to successfully park. If all the cars successfully park, we could ask what is the probability that a specific permutation $\pi\in S_{n}$ is the outcome of the parking process started with a preference list $\alpha\in[n]^{n}$.
@@
  • Other Probabilistic Parking Protocols. As mentioned in the introduction, there has been fairly extensive work studying parking on random objects such as random trees of various types, and frozen Erdős Rényi graphs. Comparatively few models of parking have been introduced which randomize the parking procedure itself, however. (For some other examples see [DamronLyu1, DamronLyu2, durmic2022probabilistic, nadeau:hal-04262134, Scott, tian2021generalizing, VarinGolf].) In many of these models, including the model discussed in this paper, the motion of the cars is characterized by random walks. In particular, the present model has deep connections to the interacting particle system IDLA. Thus it seems interesting to derive a new probabilistic parking protocol which relates to (or perhaps is independent of) other more classical particle models or stochastic processes.
@@ Appendix A Supplementary Results
$=n\left[\sum_{j=-1}^{n}j\binom{j+(n-1)-1}{j}-(n-1)\binom{2n-1}{n-1}-n\binom{2n}{n}\right]\hskip 18.06749pt\text{(Parallel summation)}$
$\hskip 36.135pt+n\binom{2n-1}{n-2}\ \ \$
@@
$\hskip 36.135pt+n\binom{2n-1}{n-2}\ \ \$
$=\frac{n^{3}(n-1)}{(n+1)(n+2)}\binom{2n-1}{n-1}$$\qed$
@@
Since the elements in $A_{n}(j)$ are uniquely determined by their first n1n-1 coordinates, and since any prefix belonging to both $B_{n}(j)$ and $B_{n}(j-t)$ for $2\leq t\leq j-1$ also belongs to $B_{n}(j-1)$, we conclude
@@
The case of $g_{0}=g_{i}=0$ is easily shown to be satisfied. For $1\leq s\leq i-1$, use that
@@
[…]
Next we evaluate the last case in (A.2) using (A.1). By (A.3), $p\frac{w_{s+1}}{w}+q\frac{w_{s-1}}{w}=1$ and so our only concern with respect to each term on the right hand side of (A.1), is the $\frac{s\left(1+\left(\frac{q}{p}\right)^{s\pm 1}\right)}{(p-q)\left(1-\left(\frac{q}{p}\right)^{s\pm 1}\right)}$ contribution. Thus, using (A.3) and (A.4), we have that
@@ Acknowledgments
The authors gratefully acknowledge that this collaboration began at the 2024 Graduate Research Workshop in Combinatorics, which is supported in part by NSF Grant DMS – 1953445. The authors would like to sincerely thank Steve Butler for initial contributions and providing a key insight in Proposition 2.2. The authors also thank Alice Contat, Ivailo Hartarsky, and Philippe Nadeau for fruitful discussion, and for introducing us to similar models in the literature. Finally, we would like to thank Persi Diaconis for taking the time to speak with us about this project, for helpful suggestions, and for offering his unique perspective on the model. Parts of this project were also presented at the Oberwolfach Mini-Workshop: Mixing Times in the Kardar-Parisi-Zhang Universality Class where useful discoveries were made, and helpful discussions were had.